OpenStudy (anonymous):

How do you rearrange this function for y in terms of x.... x=e^-7y +3 Thanks! :)

OpenStudy (zzr0ck3r):

\(x=e^{-7y}+3\) \(\large ?\)

OpenStudy (zzr0ck3r):

@VintageTypewriter ?

OpenStudy (anonymous):

Sorry...I was trying to do it :) like that

OpenStudy (zzr0ck3r):

\(x-3=e^{-7y}\implies \log(x-3)=\log(e^{-7y})\implies \log(x-3)=-7y\\y=\frac{-\log(x-3)}{7}\)

OpenStudy (zzr0ck3r):

\(\log(x)=\log_e(x) = \text{ln}(x)\)

OpenStudy (anonymous):

oh!!! Thanks :) :) Makes so much sense now!

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

so...how would you find the domain? is x not defined when log (x-3) < or equal to 0

OpenStudy (anonymous):

I mean when \[(x-3)\le0\]

OpenStudy (zzr0ck3r):

\(x-3>0\)

OpenStudy (zzr0ck3r):

you want the domain of the new function?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh yes greater than...sorry....my mistake